# On the Origin of the High Star-Formation Efficiency in Massive Galaxies at Cosmic Dawn

### v1

@@ 2.6 The suite of simulations
yesFbk $4400$ $1.0$ $3.0$ yes yes yes $3.52$ $144.3$ 19.8% $13.8\%$ $1.39$
SNeOnly $880$ $1.0$ $3.0$ yes no yes $6.03$ $293.6$ 33.9% $20.2\%$ $0.50$
photOnly $4400$ $1.0$ $3.0$ yes yes no $9.16$ $349.2$ 51.6% $17.7\%$ $0.03$
noFbk $4400$ $1.0$ $3.0$ yes no no $9.20$ $352.1$ 51.8% $18.1\%$ $0.03$
@@
highPhot $880$ $1.0$ $3.0$ yes yes yes $1.81$ $53.7$ $10.2\%$ $7.2\%$ $10.1$
medPhot $4400$ $1.0$ $3.0$ yes yes yes $3.52$ $144.3$ $19.8\%$ $13.8\%$ $1.39$
lowPhot $22000$ $1.0$ $3.0$ yes yes yes $5.06$ $275.6$ $28.5\%$ $20.7\%$ $0.56$
highPhotC $880$ $1.0$ $3.0$ no yes yes $2.05$ $62.7$ $11.6\%$ $13.8\%$ (const) $7.42$
medPhotC $4400$ $1.0$ $3.0$ no yes yes $3.12$ $121.8$ $17.6\%$ $13.8\%$ (const) $1.35$
lowPhotC $22000$ $1.0$ $3.0$ no yes yes $6.30$ $314.4$ $35.5\%$ $13.8\%$ (const) $0.57$
@@
instSNe $4400$ $1.0$ $0.0$ yes yes yes $3.58$ $142.2$ $20.2\%$ $14.4\%$ $1.14$
fastSNe $4400$ $1.0$ $0.5$ yes yes yes $3.58$ $136.4$ $20.1\%$ $14.1\%$ $1.59$
medSNe $4400$ $1.0$ $3.0$ yes yes yes $3.52$ $144.3$ $19.8\%$ $13.8\%$ $1.39$
@@
solTurb $4400$ $0.3$ $3.0$ yes yes yes $2.69$ $78.0$ $15.1\%$ $6.0\%$ $2.17$
varTurb $4400$ from $\mathbf{v}$ $3.0$ yes yes yes $3.78$ $137.8$ $21.3\%$ $12.6\%$ $2.74$
compTurb $4400$ $1.0$ $3.0$ yes yes yes $3.52$ $144.3$ $19.8\%$ $13.8\%$ $1.39$
@@
We use the early feedback series to investigate the effect of photoionization feedback in detail. In highPhot, medPhot, and lowPhot, we progressively decrease the star cluster particle mass by factors of 5. A smaller star cluster particle mass corresponds to more efficient photoionization feedback (Sec. 2.5). The fiducial star cluster particle mass in medPhot is given by the minimum baryon mass resolution $M_{\rm cl,fid}=M_{\rm b,min}=4400M_{\odot}$, which happens to be the same order as the physically-motivated value given by Equation 34.
highPhotC, medPhotC, and lowPhotC are identical to highPhot, medPhot, and lowPhot; except they do not use the MFF star formation recipe. Instead, the local SFE is set to a constant value above a density threshold, following the standard star formation recipe in galaxy simulations (e.g. Hopkins et al., 2014). The density threshold and constant local SFE are calibrated to the fiducial simulation. Specifically, the density threshold is set to the 25th-percentile star-forming density $4.24\times 10^{-21}~{}{\rm g/cm^{3}}$ and the constant local SFE is set to the median local SFE $13.8\%$.
@@ 3.2 Galaxy structure
The galaxy mass is concentrated around a single plane with a rotational velocity field, suggesting a disc structure. The disciness of the galaxy can be quantified by the scale height $H/R$ and the ratio of the rotational velocity to the radial velocity dispersion $v_{\rm rot}/\sigma_{r}$. The calculation of these metrics is described in Appendix C.
In a rotationally-supported thin disc, we expect $H/R\ll 1$ and $v_{\rm rot}/\sigma_{r}\gg 1$. For gas with temperature $T\leq 10^{4}~{}{\rm K}$, we find a disc height $\approx 326~{}{\rm pc}$ and disc radius $\approx 696~{}{\rm pc}$, giving a scale height $(H/R)_{\rm gas}\approx 0.468$. In addition, we find $v_{\rm rot}/\sigma_{r}\gg 1\approx 2.16$. For stars, we find a disc height $\approx 115~{}{\rm pc}$ and a disc radius $\approx 431~{}{\rm pc}$, giving a scale height $(H/R)_{\rm star}\approx 0.266$. In addition, we find $v_{\rm rot}/\sigma_{r}\approx 0.919$.
These numbers indicate that our galaxy is a thick, rotationally-supported gas disc surrounding a smaller stellar disc. The gas scale height calculation for the disc may be artificially inflated by accretion of cold gas. At lower redshift, MDGs may start to resemble massive star-forming galaxies at cosmic noon, which show disc structures in observation (Förster Schreiber et al., 2006; Genzel et al., 2006, 2008) and simulation (Danovich et al., 2015).
@@
At the locations of these clumps, the face-on gas surface density reaches $\Sigma_{\rm gas}\sim 10^{4}~{}M_{\odot}/{\rm pc^{2}}$ and the Hydrogen number density reaches $n_{\rm H}\sim 3\times 10^{3}~{}{\rm cm^{-3}}$. This represents a much denser environment than Milky-Way-like galaxies with face-on gas surface densities $\Sigma_{\rm gas}\sim 10-100~{}M_{\odot}/{\rm pc^{2}}$ and CNM Hydrogen number densities $n_{\rm H}\sim 10-100~{}{\rm cm^{-3}}$.
@@ 3.3 Gas lifecycle
The lifecycle of gas in our simulated galaxies can be understood in temperature-density phase space. In Figure 4, we show the distribution of gas mass in temperature-density phase space for noFbk, SNeOnly, lowPhot, medPhot, and highPhot at $z=9$. Note that the temperature is divided by $\mu$, the mean molecular weight.
We identify multiple ISM gas phases, including the hot ionized medium (HIM; $n_{\rm H}\lesssim 10~{}{\rm cm^{-3}}$ and $T\gtrsim 10^{4}~{}{\rm K}$), the warm neutral medium (WNM; $n_{\rm H}\lesssim 10~{}{\rm cm^{-3}}$ and $T\sim 10^{4}~{}{\rm K}$), the cold neutral medium (CNM; $n_{\rm H}\gtrsim 10~{}{\rm cm^{-3}}$ and $T\lesssim 10^{4}~{}{\rm K}$), and photoionized gas ($n_{\rm H}\gtrsim 10~{}{\rm cm^{-3}}$ and $T\sim 10^{4}~{}{\rm K}$). We remind the reader that the ionization state of the gas only enters into our simulations through the prescribed cooling function.
@@
The lower right panel of Figure 4 is schematic diagram illustrating the progression of gas through the phase space in steps: (i) gas from streams collapses as it accretes onto the galaxy, joining the HIM; (ii) the gas cools to the Hydrogen ionization temperature $\sim 10^{4}~{}{\rm K}$, where it becomes neutral and joins the WNM; (iii) the neutral gas cools less efficiently than the ionized gas, so it collapses while maintaining approximately the same temperature; (iv) once the density becomes sufficiently high, cooling becomes efficient and the gas joins the CNM; (v) eventually, stars begin to form and photoionization feedback generates photoionized gas at $\sim 10^{4}~{}{\rm K}$; (vi) the gas continues to collapse until it once again becomes cold and dense enough to form stars; (vii) SN explosions can recycle gas in any of the previous stages back into the WNM or HIM.
We can build up this picture by starting with our most simple simulation noFbk. In this simulation, gas accretes, cools, and collapses following steps (i)-(iii). However, without SNe to enrich the gas with metals, the gas is pristine and must reach high densities $\sim 300~{}{\rm cm^{-3}}$ before it can cool below $\sim 10^{4}~{}{\rm K}$. At this point, the gas is so dense that it only needs to cool to $\sim 1000~{}{\rm K}$ before stars form efficiently. This also describes the formation of the first generation of stars in simulations that include SNe.
The most striking difference between noFbk and SNeOnly is the presence of gas at extreme temperatures $\gtrsim 10^{7}~{}{\rm K}$. This gas must be produced by SNe exploding in diffuse gas. This can occur when a young star cluster particle diffuses into a diffuse region over its lifetime, or when the local gas conditions around a star cluster particle are modified by early feedback or other SNe.
@@ 3.4 Local gas properties
In Figures 5 and 6, we show the mass-weighted and volume-weighted distributions respectively of gas density, temperature, turbulent velocity dispersion, and metallicity for the early feedback and turbulence forcing series. We plot turbulent velocity dispersion rather than turbulent Mach number to isolate the differences in turbulence between simulations from the temperature-dependence $\mathcal{M}_{\rm turb}\propto T^{-0.5}$ at a constant TKE.
@@
HII regions created by photoionization feedback produce a sharp peak in the temperature distribution at $\sim 10^{4}~{}{\rm K}$. The density distribution does not change significantly across the early feedback series, suggesting that the thermal pressure from photoionization feedback is insufficient to halt the collapse of the star-forming clouds. This is expected, because for high mass star-forming clouds $M_{\rm cl}\gtrsim 10^{6}~{}M_{\odot}$, the sound speed $c_{\rm s}\sim 10~{}{\rm km/s}$ in photoionized gas is less than the escape speed of the cloud (e.g. Krumholz & Matzner, 2009).
The volume-weighted distributions tell a similar story. In this case, the peak in the temperature distribution at $T\sim 10^{4}~{}{\rm K}$ represents the volume of WNM rather photoionized gas, whose volume fraction is generally smaller. As photoionization feedback becomes more efficient, volume moves from ultra-diffuse HIM gas $\lesssim 10^{-2}~{}{\rm cm^{-3}}$ to WNM gas $\sim 10^{-1}~{}{\rm cm^{-3}}$.
Solenoidal forcing decreases the local SFE for the same gas conditions and TKE. Therefore, gas must reach higher densities to form stars efficiently, resulting in more gas mass at high densities $\gtrsim 10^{3}~{}{\rm cm^{-3}}$.
Figure 5: Histograms of the hydrogen number density (first column), modified temperature (second column), turbulent velocity dispersion (third column), and metallicity (fourth column) in a box of sidelength $5~{}{\rm kpc}$ around the most massive galaxy at $z=9$ in the early feedback series (top row) and turbulence forcing series (bottom row), weighted by gas mass. Each simulation is shown in a different color, with the fiducial simulation in blue. An additional histogram of the turbulence forcing parameter is included for the turbulence forcing simulation.
Figure 6: Histograms of the hydrogen number density (first column), modified temperature (second column), turbulent velocity dispersion (third column), and metallicity (fourth column) in a box of sidelength $5~{}{\rm kpc}$ around the most massive galaxy at $z=9$ in the early feedback series (top row) and turbulence forcing series (bottom row), weighted by volume. Each simulation is shown in a different color, with the fiducial simulation in blue. An additional histogram of the turbulence forcing parameter is included for the turbulence forcing simulation.
@@
Figure 7: Histograms of the hydrogen number density (first column), modified temperature (second column), turbulent velocity dispersion (third column), and metallicity (fourth column) of star-forming cells in the early feedback series (top row) and turbulence forcing series (bottom row), weighted by stellar mass. Each simulation is shown in a different color, with the fiducial simulation in blue. The inset plot shows a histogram of the turbulence forcing parameter for the turbulence forcing series. An additional histogram of the turbulence forcing parameter is included for the turbulence forcing simulation.
@@
Figure 8: Histograms of the hydrogen number density (first column), modified temperature (second column), turbulent velocity dispersion (third column), and metallicity (fourth column) of cells with SN events in the early feedback series (top row) and turbulence forcing series (bottom row), weighted by SN number. Each simulation is shown in a different color, with the fiducial simulation in blue. An additional histogram of the turbulence forcing parameter is included for the turbulence forcing simulation.
@@
Our finding contradicts the prediction of the FFB model (Dekel et al., 2023) that the SN delay time provides a crucial window of opportunity for efficient star formation. However, our ability to test the FFB model is limited by our resolution $\Delta x_{\rm min}\simeq 10~{}{\rm pc}$. We cannot fully resolve the parsec-scale dynamics which control star-forming clouds, especially in dense regions where the SN cooling radius is unresolved (Sec. 4.7).

### v2

@@ 2.6 The suite of simulations
yesFbk $4400$ $1.0$ $3.0$ yes yes yes $3.52$ $46.7$ 19.8% $13.8\%$ $1.39$
SNeOnly $880$ $1.0$ $3.0$ yes no yes $6.03$ $51.5$ 33.9% $20.2\%$ $0.50$
photOnly $4400$ $1.0$ $3.0$ yes yes no $9.16$ $82.5$ 51.6% $17.7\%$ $0.03$
noFbk $4400$ $1.0$ $3.0$ yes no no $9.20$ $83.2$ 51.8% $18.1\%$ $0.03$
@@
highPhot $880$ $1.0$ $3.0$ yes yes yes $1.81$ $17.3$ $10.2\%$ $7.2\%$ $10.1$
medPhot $4400$ $1.0$ $3.0$ yes yes yes $3.52$ $46.7$ $19.8\%$ $13.8\%$ $1.39$
lowPhot $22000$ $1.0$ $3.0$ yes yes yes $5.06$ $56.0$ $28.5\%$ $20.7\%$ $0.56$
highPhotC $880$ $1.0$ $3.0$ no yes yes $2.05$ $23.0$ $11.6\%$ $13.8\%$ (const) $7.42$
medPhotC $4400$ $1.0$ $3.0$ no yes yes $3.12$ $41.8$ $17.6\%$ $13.8\%$ (const) $1.35$
lowPhotC $22000$ $1.0$ $3.0$ no yes yes $6.30$ $66.5$ $35.5\%$ $13.8\%$ (const) $0.57$
@@
instSNe $4400$ $1.0$ $0.0$ yes yes yes $3.58$ $43.4$ $20.2\%$ $14.4\%$ $1.14$
fastSNe $4400$ $1.0$ $0.5$ yes yes yes $3.58$ $40.9$ $20.1\%$ $14.1\%$ $1.59$
medSNe $4400$ $1.0$ $3.0$ yes yes yes $3.52$ $46.7$ $19.8\%$ $13.8\%$ $1.39$
@@
solTurb $4400$ $0.3$ $3.0$ yes yes yes $2.69$ $32.9$ $15.1\%$ $6.0\%$ $2.17$
varTurb $4400$ from $\mathbf{v}$ $3.0$ yes yes yes $3.78$ $56.4$ $21.3\%$ $12.6\%$ $2.74$
compTurb $4400$ $1.0$ $3.0$ yes yes yes $3.52$ $46.7$ $19.8\%$ $13.8\%$ $1.39$
@@
We use the early feedback series to investigate the effect of photoionization feedback in detail. In highPhot, medPhot, and lowPhot, we progressively decrease the star particle mass by factors of 5. A smaller star particle mass corresponds to more efficient photoionization feedback (Sec. 2.5). The fiducial star cluster particle mass in medPhot is given by the minimum baryon mass resolution $M_{\rm cl,fid}=M_{\rm b,min}=4400M_{\odot}$ close to the physically-motivated value given by Equation 34.
highPhotC, medPhotC, and lowPhotC are identical to highPhot, medPhot, and lowPhot, except they do not use the MFF star formation recipe. Instead, the local SFE is set to a constant value above a density threshold, following the standard star formation recipe in galaxy simulations (e.g. Hopkins et al., 2014). The density threshold and constant local SFE are calibrated to the fiducial simulation. Specifically, the density threshold is set to the 25th-percentile star-forming density $4.24\times 10^{-21}~{}{\rm g/cm^{3}}$ and the constant local SFE is set to the median local SFE $13.8\%$.
@@ 3.2 Galaxy structure
The galaxy mass is concentrated around a single plane with a rotational velocity field, suggesting a disc structure. Our estimate of the galaxy radius $720~{}{\rm pc}$ in Appendix A agrees nicely with Figure 3. The morphology of the galaxy can be quantified by the scale height $H/R$ and the ratio of the rotational velocity to the radial velocity dispersion $v_{\rm rot}/\sigma_{r}$. The calculation of these metrics is described in Appendix C.
In a rotationally-supported thin disc, we expect $H/R\ll 1$ and $v_{\rm rot}/\sigma_{r}\gg 1$. The relative contribution of the centrifugal force to the support is (Ceverino et al., 2012)
$\mathcal{R}=\left[1+2(\sigma_{r}/v_{\rm rot})^{2}\right]^{-1}$ (36)
For gas with temperature $\leq 10^{4}~{}{\rm K}$, we find a disc height $\approx 326~{}{\rm pc}$ and disc radius $\approx 696~{}{\rm pc}$, giving a scale height $(H/R)_{\rm gas}\approx 0.468$, although this may be artificially inflated by accreting cold gas in the polar regions. In addition, we find $v_{\rm rot}/\sigma_{r}\approx 2.16$, corresponding to $\mathcal{R}\simeq 0.70$. This implies that the gas disc is mostly supported by rotation (70%) with some turbulent support (30%).
Likewise, for stars, we find a disc height $\approx 115~{}{\rm pc}$ and a disc radius $\approx 431~{}{\rm pc}$, giving a scale height $(H/R)_{\rm star}\approx 0.266$. In addition, we find $v_{\rm rot}/\sigma_{r}\approx 0.919$, corresponding to $\mathcal{R}\simeq 0.30$. This implies that the stellar disc is mostly supported by turbulence (70%) with some rotational support (30%). The significant contributions from both rotation and turbulence explain the thick disk structure. At lower redshift, MDGs may start to resemble massive star-forming galaxies at cosmic noon, which show disc structures in observation (Förster Schreiber et al., 2006; Genzel et al., 2006, 2008) and simulation (Danovich et al., 2015).
@@
The average face-on baryonic surface density within $1~{}{\rm kpc}$ of the center is $1500~{}M_{\odot}/{\rm pc^{2}}$, with a contribution $470~{}M_{\odot}/{\rm pc^{2}}$ from the gas. This is less than our estimate $\sim 10^{4}~{}M_{\odot}/{\rm pc^{2}}$ of the baryonic surface density in Appendix A. The discrepancy is primarily due to baryons inside the halo which have not yet settled into the disk, making $f_{\rm b}M_{\rm vir}$ an overestimate of the baryon mass in the disk. The baryon mass inside the galaxy radius $720~{}{\rm pc}$ is only a third of the baryon mass inside the virial radius $14~{}{\rm kpc}$.
The surface density is higher at smaller radii, but the disc is increasingly dominated by stars. The face-on baryonic surface density within $180~{}{\rm pc}$ of the center is $8000~{}M_{\odot}/{\rm pc^{2}}$ with a contribution $880~{}M_{\odot}/{\rm pc^{2}}$ from the gas. Gas density is enhanced significantly within the clumps, with face-on gas surface density $\sim 10^{4}~{}M_{\odot}/{\rm pc^{2}}$ and Hydrogen number density $\sim 3\times 10^{3}~{}{\rm cm^{-3}}$. Most of the star formation occurs within these clumps.
@@ 3.3 Gas lifecycle
The lifecycle of gas in our simulated galaxies can be understood in temperature-density phase space. In Figure 4, we show the distribution of gas mass in temperature-density phase space for noFbk, SNeOnly, lowPhot, medPhot, and highPhot at $z=9$. The temperature is divided by $\mu$, the mean molecular weight.
We identify multiple ISM gas phases, including the hot ionized medium (HIM; $n_{\rm H}\sim 0.005~{}{\rm cm^{-3}}$ and $T\gtrsim 10^{5.5}~{}{\rm K}$), the warm neutral medium (WNM; $n_{\rm H}\sim 0.6~{}{\rm cm^{-3}}$ and $T\sim 5000~{}{\rm K}$), the cold neutral medium (CNM; $n_{\rm H}\sim 30~{}{\rm cm^{-3}}$ and $T\sim 100~{}{\rm K}$), and photoionized gas ($n_{\rm H}\gtrsim 0.3~{}{\rm cm^{-3}}$ and $T\sim 10^{4}~{}{\rm K}$). We remind the reader that the ionization state of the gas only enters into our simulations through the prescribed cooling function.
@@
The gas mass in the galaxy increases over time due to accretion from the IGM. If the cooling time near the virial radius is short, this precludes the formation of a stable accretion shock, and the cold gas accretes directly onto the central galaxy (cold mode accretion). Otherwise, the shock-heated gas must cool in the circum-galactic medium (CGM) before reaching the galaxy (hot mode accretion) (Birnboim & Dekel, 2003; Kereš et al., 2005). Figure 4 probes a region deep inside the virial radius, so it does not explicitly separate these two scenarios.
Hot mode accretion typically dominates at halo masses $\gtrsim 10^{11}$ to $10^{12}~{}M_{\odot}$ for (quasi-)spherical accretion (Birnboim & Dekel, 2003; Kereš et al., 2005). However, at high redshift, accretion in massive haloes is increasingly dominated by dense filaments, shortening the cooling time and enhancing cold mode accretion (Dekel & Birnboim, 2006; Kereš et al., 2009). Therefore, we expect MDGs to accrete primarily in the cold mode despite the large halo mass and low metallicity of the accreting gas. The CGM is poorly resolved in our simulations, so we leave a detailed study of gas accretion onto MDGs to future work.
The lower right panel of Figure 4 is schematic diagram illustrating the processes that move gas through the phase space: (i) cold gas from cosmic filaments accrete onto the galaxy; (ii) some of the gas is shock-heated by turbulence, joining the HIM; (iii) HIM gas cools to the Hydrogen ionization temperature $\sim 10^{4}~{}{\rm K}$, where it becomes neutral and joins the WNM; (iv) the neutral gas cools less efficiently than the ionized gas, so it collapses while maintaining approximately the same temperature; (v) once the density becomes sufficiently high, cooling becomes efficient and the gas joins the CNM; (vi) eventually, stars begin to form and photoionization feedback generates photoionized gas at $\sim 10^{4}~{}{\rm K}$; (vii) the gas continues to collapse until it once again becomes cold and dense enough to form stars; (viii) SN explosions can recycle gas in any of the previous stages back into the WNM or HIM.
We can build up this picture by starting with our most simple noFbk simulation. In this simulation, gas accretes, cools, and collapses following steps (i)-(iv). However, without SNe to enrich the gas with metals, the gas is pristine and must reach high densities $\sim 300~{}{\rm cm^{-3}}$ before it can cool below $\sim 10^{4}~{}{\rm K}$. At this point, the gas is so dense that it only needs to cool to $\sim 1000~{}{\rm K}$ before stars form efficiently. This also describes the formation of the first generation of stars in the other simulations that include SNe.
The most striking difference between noFbk and SNeOnly is the presence of gas at temperatures $\gtrsim 10^{7}~{}{\rm K}$. Only SNe can produce gas at these extreme temperatures (McCray & Snow, 1979; Spitzer, 1990). The virial temperature reached by the accretion shock in hot mode accretion (Eq. 55) and the temperatures reached by turbulent shocks (Wada & Norman, 2001; Ceverino & Klypin, 2009) are both $\sim 10^{6}~{}{\rm K}$. If a SN explodes in diffuse gas, the ejecta reaches a far greater temperature given by
$\frac{T_{\rm SN}}{\mu}=\frac{m_{\rm p}}{k_{\rm B}}\frac{E_{\rm SN}}{m_{\rm ej}}\simeq 6.1\times 10^{8}E_{51}m_{{\rm ej},10}~{}{\rm K}$ (37)
where $E_{51}=E_{\rm SN}/10^{51}~{}{\rm erg}$ and $m_{\rm ej,10}=m_{\rm ej}/10~{}M_{\odot}$ are the SN energy and ejecta mass respectively. SN can explode in diffuse gas when a young star particle diffuses into a diffuse region over its lifetime, or when the local gas conditions around a star cluster particle are modified by early feedback or other SNe.
@@ 3.4 Local gas properties
In Figures 5 and 6, we show the mass-weighted and volume-weighted distributions respectively of gas density, temperature, turbulent velocity dispersion, and metallicity for the early feedback and turbulence forcing series. We plot turbulent velocity dispersion $\sigma_{\rm turb}^{2}=(2/3)\varepsilon_{\rm turb}/\rho$ rather than turbulent Mach number to isolate the differences in turbulence between simulations from the temperature-dependence $\mathcal{M}_{\rm turb}\propto T^{-0.5}$ at a constant TKE.
@@
HII regions created by photoionization feedback produce a sharp peak in the temperature distribution at $\sim 10^{4}~{}{\rm K}$. The density distribution does not change significantly across the early feedback series, suggesting that the thermal pressure from photoionization feedback is insufficient to halt the collapse of the star-forming clouds. This is expected, because for high mass star-forming clouds $\gtrsim 10^{6}~{}M_{\odot}$, the sound speed $\sim 10~{}{\rm km/s}$ in photoionized gas is less than the escape speed of the cloud (e.g. Krumholz & Matzner, 2009).
The volume-weighted distributions tell a similar story. In this case, the peak in the temperature distribution at $T\sim 10^{4}~{}{\rm K}$ represents the volume of WNM rather photoionized gas, whose volume fraction is generally smaller. As photoionization feedback becomes more efficient, volume moves from HIM gas to less diffuse WNM gas.
Solenoidal forcing decreases the local SFE for the same gas conditions and TKE. Therefore, gas must reach higher densities to form stars efficiently, resulting in more gas mass at high densities in solTurb.
Figure 5: Histograms of the hydrogen number density (first column), modified temperature (second column), turbulent velocity dispersion (third column), and metallicity (fourth column) in a box of sidelength $5~{}{\rm kpc}$ around the most massive galaxy at $z=9$ in the early feedback series (top row) and turbulence forcing series (bottom row), weighted by gas mass. Each simulation is shown in a different color, with the fiducial simulation in blue. An additional histogram of the turbulence forcing parameter is included for the variable turbulence forcing simulation.
Figure 6: Histograms of the hydrogen number density (first column), modified temperature (second column), turbulent velocity dispersion (third column), and metallicity (fourth column) in a box of sidelength $5~{}{\rm kpc}$ around the most massive galaxy at $z=9$ in the early feedback series (top row) and turbulence forcing series (bottom row), weighted by volume. Each simulation is shown in a different color, with the fiducial simulation in blue. An additional histogram of the turbulence forcing parameter is included for the variable turbulence forcing simulation.
@@
Figure 7: Histograms of the hydrogen number density (first column), modified temperature (second column), turbulent velocity dispersion (third column), and metallicity (fourth column) of star-forming cells in the early feedback series (top row) and turbulence forcing series (bottom row), weighted by stellar mass. Each simulation is shown in a different color, with the fiducial simulation in blue. The inset plot shows a histogram of the turbulence forcing parameter for the turbulence forcing series. An additional histogram of the turbulence forcing parameter is included for the variable turbulence forcing simulation.
@@
Figure 8: Histograms of the hydrogen number density (first column), modified temperature (second column), turbulent velocity dispersion (third column), and metallicity (fourth column) of cells with SN events in the early feedback series (top row) and turbulence forcing series (bottom row), weighted by SN number. Each simulation is shown in a different color, with the fiducial simulation in blue. An additional histogram of the turbulence forcing parameter is included for the variable turbulence forcing simulation.
@@
Our finding contradicts the prediction of the FFB model (Dekel et al., 2023) that the SN delay time provides a crucial window of opportunity for efficient star formation. However, our ability to test the FFB model is limited by our effective resolution $\simeq 10~{}{\rm pc}$. We cannot fully resolve the parsec-scale dynamics which control star-forming clouds, especially in dense regions where the SN cooling radius is unresolved (Sec. 4.7).